# Causal Inference with Large Language Model: A Survey

### v1

@@ 4 Evaluations of LLMs in Causal Tasks
4.1 Overview
In this section, we summarize recent evaluation results of LLMs in causal tasks. We mainly focus on causal discovery and causal effect estimation, and also introduce several representative tasks spanning Rung 1$\sim$3. A collection of datasets used in LLM-related causal tasks is shown in Table 1. In Table 2 and Table 3, we compare the performance of different LLMs in different tasks (including causal discovery and other tasks spanning different rungs in causal ladder) on multiple datasets. The mentioned LLMs include ada, babbage, curie, davinci Brown et al. (2020), text-ada-001, text-babbage-001, text-curie-001, text-davinci-001, text-davinci-002, text-davinci-003 Ouyang et al. (2022), Llama 2 (7B, 13B, 70B) Touvron et al. (2023), and OpenAI’s GPT series Achiam et al. (2023); OpenAI (2022).
@@ 4.2 LLM for Causal Discovery
Table 2: Performance (accuracy) of different models in causal discovery tasks on different datasets, including CausalEffectPairs (CEpairs for short), E-CARE, COPA, CALM-CA, and Neuro Pain. In the columns in white (CausalEffectPairs, E-CARE, COPA), the models are evaluated for the pairwise causal discovery task; In the column in gray, the models are evaluated for the causal attribution task; in the column in cyan, the models are evaluated for the full graph discovery task. In the upper part, we show results with basic prompt; while in the lower part, we show results of GPT-4 with different prompting strategies. We also present results under prompts in the form of binary "yes/no" questions and multi-choice questions. The results are collected from Kıcıman et al. (2023) and Chen et al. (2024a). Note that the experimental settings such as prompt templates may be different.
@@
Table 3: Performance (accuracy) of different models in causal tasks in the ladder of causation (Rung 1 $\sim$ Rung 3) on different datasets, including CLADDER, CaLM, CRASS, and E-CARE. The column in gray correspond to tasks in Rung 1 (corr=correlation), the columns in white involve tasks in Rung 2 (ATE=average treatment effect; CDE = controlled direct effect; BAJ= backdoor adjustment); the columns in cyan correspond to tasks in Rung 3 (CR=counterfactual reasoning; NDE=natural direct effect; CE=causal explanation). In the upper part, we show results with the basic prompt; while in the lower part, we show results of GPT-4 with different prompting strategies. The results are collected from Chen et al. (2024a) and Jin et al. (2023a). Note that the experimental settings such as prompt templates may be different.
Causal discovery aims to identify the causal relationships between different variables, often serving as a fundamental step in real-world causal analysis. Most traditional causal discovery approaches rely on the data values and use statistical approaches to infer the underlying causal structure over the corresponding variables. These approaches include constraint-based methods (e.g., PC algorithm Spirtes et al. (2000) and FCI algorithm Spirtes et al. (2013); Zhang (2008)) which infer causal relationships by leveraging conditional independence tests, and score-based methods which assign scores to candidate causal graphs w.r.t. certain scoring criterion and seek the candidate causal graph with the highest score (e.g., GES algorithm Chickering (2002)). Various classical statistical approaches and recent machine learning or deep learning technologies Spirtes and Zhang (2016); Nogueira et al. (2022); Vowels et al. (2022) have been used in causal discovery.
Recent developments in LLMs provide new perspectives for causal discovery Takayama et al. (2024); Liu et al. (2024a); Kıcıman et al. (2023). Different from most existing causal discovery methods which can only utilize the data values of variables, LLMs can also leverage the metadata (e.g., the names of variables, the problem context) related to these variables to discover the implicit causal relationships. This reasoning process makes LLM-based causal discovery closer to human recognition. Recent literature Kıcıman et al. (2023) refer to this ability as knowledge-based causal discovery, and their experiments show that LLM-based knowledge-based causal discovery outperforms existing causal discovery methods on benchmarks Mooij et al. (2016). Currently, a variety of investigations have been conducted on LLMs in causal discovery tasks Kıcıman et al. (2023); Cai et al. (2023); Gao et al. (2023); Jin et al. (2023b); Long et al. (2023). These investigations are often conducted in the form of multi-choice or free-text question-answering, and they can mainly be divided into two types: pairwise causal discovery and full causal graph discovery.
Pairwise causal discovery (PCD) focuses on a pair of variables, either aiming to infer the causal direction ($A\rightarrow B$ or $A\leftarrow B$) between a given pair of variables $(A,B)$, or aiming to judge the existence of a causal relation between two variables. Among them, the experiments in Kıcıman et al. (2023) use the names of variables when constructing prompts, and their results show that LLMs (including GPT-3.5 and GPT-4) outperform state-of-the-art methods both on datasets with common variables (e.g., CauseEffectPairs Mooij et al. (2016)) and datasets that require particular domain knowledge (e.g., neuropathic pain Tu et al. (2019)). Despite the encouraging results, the empirical analysis from Zečević et al. (2023) implies that in many cases, LLMs are just “causal parrots" that repeat the embedded causal knowledge. A comparison between ChatGPT and fine-tuned small pre-trained language models Gao et al. (2023) shows LLMs’ advantage in some causal discovery tasks, but this work also discusses that the ability of LLMs in determining the existence of a causal relationship is worse than simply selecting the cause or effect of an input event from given options. Jin et al. (2023b) proposes a correlation-to-causation inference (Corr2Cause) task to evaluate the causal inference performance of LLMs. Their experimental results reveal that LLMs perform almost close to random on the task, even though this issue could be mitigated through fine-tuning, these models still have limitations in generalization on out-of-distribution settings.
@@
Full causal graph discovery aims to identify the full causal graph that describes the causal relationships among a given set of variables. Compared with pairwise causal discovery, discovering the full causal graph is a more complicated problem as it involves more variables. In a preliminary exploration Long et al. (2023), GPT-3 shows good performance in discovering the causal graph with 3-4 nodes for well-known causal relationships in the medical domain. In more complicated scenarios, the ability of different versions of GPT to discover causal edges Kıcıman et al. (2023) has been validated on the neuropathic pain dataset Tu et al. (2019) with 100 pairs of true/false causal relations. LLM-based discovery (GPT-3.5 and GPT-4) on Arctic sea ice dataset Huang et al. (2021) has comparable or even better performance than representative baselines including NOTEARS Zheng et al. (2018) and DAG-GNN Yu et al. (2019). In Ban et al. (2023), the combination of the causal knowledge generated by LLMs and data-driven methods brings improvement in causal discovery in data from eight different domains with small causal graphs (5$\sim$48 variables and 4$\sim$84 causal relations). But similarly to PCD, LLMs also face many doubts and debates about their true ability in full graph discovery.
@@ 4.3 LLM for Causal Effect Estimation
Causal effect estimation aims to quantify how much manipulating a treatment can causally influence an outcome. In most cases, the causal effect of interest is estimated from observational data. Researchers in the NLP community have also made lots of efforts in causal effect estimation from text data Feder et al. (2022); Keith et al. (2020). Causal effect estimation on text data faces unique challenges due to the high-dimensional and complicated nature, for example, some important assumptions (e.g., positivity assumption D’Amour et al. (2021)) in traditional causal effect estimation are easily violated when high-dimensional text information is a confounder Keith et al. (2020). Fortunately, the NLP progress in recent decades, such as word embeddings Almeida and Xexéo (2019), topic modeling Blei et al. (2003) and dependency parsing Nivre (2005) have significantly contributed to estimating causal effects on text.
LLM has shown impressive performance in causal effect estimation as well. Recently, the connection between causal effect estimation and LLMs includes two different branches: (1) Causal effect in data: In this setting, LLMs aim to estimate the causal effect inside data Lin et al. (2023); Kıcıman et al. (2023) by leveraging their reasoning capability and properties (e.g., ability to handle large-scale training corpora). A benchmark for the capability of LLMs in causal inference, CLADDER Jin et al. (2023a), includes query types regarding causal effect estimation at different levels, e.g., ATE, ATT, NDE, NIE in the Rung 2, and ATT, NDE, NIE in the Rung 3. Existing evaluations show that the causal effect estimation task is still quite challenging for most LLMs. However, an encouraging finding is that proper techniques such as chain-of-thought (CoT) prompting strategy Jin et al. (2023a) can improve the performance significantly. (2) Causal effect in model: This setting aims to analyze the causal effect that involves the LLM itself. Most commonly, we focus on the causal effect of input data, model neurons, or learning strategies on LLMs’ predictions Vig et al. (2020); Meng et al. (2022); Stolfo et al. (2022). These studies can reveal the underlying LLM behavior and promote further investigations such as bias elimination Vig et al. (2020), model editing Meng et al. (2022), and robustness quantification Stolfo et al. (2022). For example, Stolfo et al. (2022) explores the causal effect of input (e.g., problem description and math operators) on output solutions in LLM-based mathematical reasoning. In Vig et al. (2020), a causal mediation analysis for gender bias propagated from model input to output is conducted in language models.
@@ 4.4 LLM for Other Causal Tasks
Experiments Chen et al. (2024a); Jin et al. (2023a); Kıcıman et al. (2023) have shown that there are various other causal inference tasks that LLMs can bring benefits to. (1) Causal attribution: LLMs show their capability in attribution tasks Kıcıman et al. (2023); Cai et al. (2023) typically in the forms of "why" or "what is the cause" questions. Related tasks also include identifying necessary or sufficient causes Liu et al. (2023); Kıcıman et al. (2023). By embedding human knowledge and cultural common sense, the results show that LLMs have the potential to flexibly address attribution problems in specific domains (such as law, economics, and medicine) where conventional methods may fall short Kıcıman et al. (2023). (2) Counterfactual reasoning: Recent studies Kıcıman et al. (2023); Jin et al. (2023a) conduct experiments on LLMs in different counterfactual reasoning scenarios, which are often in "what if" questions. While this task is one of the most challenging tasks in causal inference, the demonstrated improvement in LLMs compared to other methods is noteworthy. (3) Causal explanation: Many recent works investigate causal explanations based on queries on LLMs Bhattacharjee et al. (2023); Gat et al. (2023); Cai et al. (2023); Gao et al. (2023). Despite ongoing debates regarding LLM’s actual ability for causal reasoning, most empirical studies positively indicate that LLMs serve as effective causal explainers Gao et al. (2023). Such achievement is powered by LLMs’ capability of analyzing language logic and responding to questions using natural language.
@@ 4.5 Main Observations and Insights
From the evaluation discussed above and results shown in Table 2 and Table 3, we summarzie the main observations as follows: (1) Model performance: In general, many LLMs exhibit impressive performance in various causal tasks, especially in causal discovery, even with basic prompts. In some cases, their performance can be comparable to or even surpass human-level reasoning Kıcıman et al. (2023). However, as the task difficulty increases from Rung 1 to Rung 3, their performance becomes less satisfactory in higher-level complicated causal reasoning tasks Chen et al. (2024a). (2) Enhancement through proper strategies: The performance of LLMs can be significantly enhanced with effective prompting strategies, such as few-shot ICL and CoT. These approaches enable models to leverage contextual information and reasoning processes to achieve better results. Additionally, these models can provide valuable insights through causal explanations. (3) General patterns: While no definitive laws determine model performance universally, certain trends are still observable. For instance, scaling laws suggest that larger models generally perform better, although this is not always that straightforward. These trends provide valuable insights that can guide the future design and development of models. (4) Variability in model effectiveness: There is currently no universally superior LLM or strategy for causal tasks, as their effectiveness can vary significantly depending on the specific scenario. These observations highlight the need for more nuanced and adaptable approaches. (5) Common issues in LLMs for causal tasks: Current LLMs still struggle with many issues in causal tasks. For example, the answers often lack robustness and are highly sensitive to changes in prompts Kıcıman et al. (2023); Jin et al. (2023a). Besides, these models frequently default to memorizing and repeating information rather than engaging in actual causal reasoning Zečević et al. (2023), which can limit their effectiveness in complex causal scenarios. Besides, LLMs often fail to generate self-consistent answers for causal queries, i.e., the answers from LLMs often present causal relationships that conflict with each other. Ongoing debates and criticisms about whether LLM truly performs causal inference also compel more in-depth and precise analysis and evaluation.
@@ 5 Discussion and Future Prospects
In general, LLMs offer intriguing perspectives on causal inference, but current research also reveals many limitations, pointing to potential directions for future work that could advance the field Zhang et al. (2023); Kıcıman et al. (2023). Here, we outline several promising avenues for exploration in this research area, including:
  • Incorporating human knowledge: Many causal inference tasks necessitate human knowledge. A more comprehensive and intelligent integration of human knowledge (e.g., domain expertise) into LLMs could enhance causal reasoning, enabling interdisciplinary causal knowledge integration and analysis in both general settings and specialized domains such as finance, healthcare, and law Chen et al. (2024b).
  • Improving data generation: Real-world data often lack verified causal relations and counterfactuals. Utilizing LLMs for causal data generation can provide more diverse and realistic data, enriching real-world datasets with reliable causal relationships and improving training for causal reasoning models.
  • Addressing hallucinations: Hallucinations widely exist in LLM-generated causal inference answers. Focused efforts to eliminate hallucinations can lead to more accurate and reliable causal reasoning. Additionally, addressing broader issues such as fairness and biases can further enhance the trustworthiness of LLM in causal reasoning tasks.
  • Improving explanation and interactivity: Developing more interpretable and instructable LLMs in causal reasoning is crucial. Methods such as fine-tuning, model distillation, probing, and prompt engineering can improve how LLMs interact with users. Optimizing reasoning chains to accommodate human instructions and feedback can also foster more collaborative and controllable causal inference between humans and AI.
  • Exploring multimodal causality: Real-world scenarios often involve multiple modalities. Recent studies have begun exploring causality across different modalities, such as images Li et al. (2024) and videos Lam et al. (2024). Future research could further investigate these multimodal approaches to enhance causal understanding.
  • Developing a unified causal benchmark: There is currently no unified and widely recognized benchmark for evaluating causal performance in LLMs. Creating a comprehensive benchmark would facilitate better comparison and evaluation of different models, leading to more consistent and reliable assessments.
  • Addressing limitations in causal tasks and datasets: Existing causal tasks and benchmark datasets are relatively simplistic, often limited to multiple-choice and binary classification formats. These tasks frequently lack the complexity needed to address intricate scenarios and domain-specific contexts.
  • Advancing causality-focused models: Most existing methods only evaluate original LLMs without sufficient focus on causality-centric model designs. There is a significant opportunity for further research and development in this area to deepen models’ understanding of causality and improve their effectiveness in causal inference tasks.
@@ 6 Limitations
In this survey paper, it is important to acknowledge certain limitations that shape the scope and focus of our review. Firstly, our analysis is primarily centered on the application of large language models (LLMs) for causal inference tasks, thereby excluding exploration into how causality is utilized within LLM frameworks themselves. This decision provides a targeted perspective on leveraging LLMs to enhance causal inference methodologies but does not delve into the internal mechanisms or implementations of causal reasoning within these models.
Secondly, while we comprehensively examine the technical aspects and methodological advancements in using LLMs for causal inference, we do not extensively discuss ethical considerations or potential societal impacts associated with these applications. Ethical dimensions, such as fairness, bias mitigation, and privacy concerns, are critical in the deployment of AI technologies, including LLMs, and warrant dedicated attention and scrutiny in future research and applications. Addressing these limitations ensures a nuanced understanding of the opportunities and challenges in harnessing LLMs for causal inference while also advocating for responsible and ethical AI development and deployment practices.

### v2

@@ 4 Evaluations of LLMs in Causal Tasks
This section summarizes recent evaluation results of LLMs in causal tasks. We mainly focus on causal discovery and causal effect estimation, and also introduce several representative tasks spanning Rung 1$\sim$3. A collection of datasets used in LLM-related causal tasks is shown in Table 1. In Table 2 and Table 3, we compare the performance of different LLMs in different tasks (including causal discovery and other tasks spanning different rungs) on multiple datasets. The mentioned LLMs include ada, babbage, curie, davinci Brown et al. (2020), text-ada-001, text-babbage-001, text-curie-001, text-davinci-001, text-davinci-002, text-davinci-003 Ouyang et al. (2022), Llama 2 (7B, 13B, 70B) Touvron et al. (2023), and OpenAI’s GPT series Achiam et al. (2023); OpenAI (2022).
@@ 4.1 LLM for Causal Discovery
Table 2: Performance (accuracy) of LLMs for causal discovery. Datasets include CausalEffectPairs (CEpairs), E-CARE, COPA, CALM-CA, and Neuro Pain. The white columns evaluate LLMs on PCD; in the gray column, on causal attribution; and in the cyan column, on full graph discovery. The upper part shows results with basic prompts; the lower part shows GPT-4 results with different prompting strategies. The tasks are shown in either binary "yes/no" or multi-choice formats. Results are drawn from Kıcıman et al. (2023) and Chen et al. (2024a).
@@
Table 3: Performance (accuracy) of LLMs in causal tasks across the ladder of causation (Rung 1$\sim$3) on datasets including CLADDER, CaLM, CRASS, and E-CARE. The gray column shows results for Rung 1 (corr=correlation), the white columns for Rung 2 (ATE=average treatment effect; CDE = controlled direct effect; BAJ= backdoor adjustment), and the cyan columns for Rung 3 (CR=counterfactual reasoning; NDE=natural direct effect; CE=causal explanation). The upper part shows results with basic prompts, while the lower part presents GPT-4 results with different prompting strategies. Data is sourced from Chen et al. (2024a) and Jin et al. (2023a).
Unlike traditional causal discovery methods that rely solely on data values Spirtes et al. (2000, 2013); Chickering (2002), LLMs can also leverage metadata (e.g., variable names, problem context) to uncover implicit causal relationships, making their reasoning closer to human cognition. Many studies have explored LLMs for causal discovery Kıcıman et al. (2023); Cai et al. (2023); Gao et al. (2023); Jin et al. (2023b); Long et al. (2023), focusing on pairwise causal discovery and full causal graph discovery, often framed as multi-choice or free-text question-answering tasks.
Pairwise causal discovery (PCD). PCD focuses on inferring the causal direction ($A\rightarrow B$ or $A\leftarrow B$) or determining the existence of a causal relationship. Kıcıman et al. (2023) use variable names in prompts, showing that LLMs (e.g., GPT-3.5, GPT-4) outperform state-of-the-art methods on datasets like CauseEffectPairs Mooij et al. (2016) and domain-specific datasets like neuropathic pain Tu et al. (2019). However, some studies Zečević et al. (2023) suggest LLMs often act as "causal parrots", merely repeating embedded causal knowledge. Jin et al. (2023b) proposes a correlation-to-causation inference (Corr2Cause) task, where LLMs performed close to random. Although fine-tuning improved their performance, they still struggle with generalization in out-of-distribution scenarios. In summary, many studies Gao et al. (2023); Kıcıman et al. (2023); Jin et al. (2023b, a); Chen et al. (2024a) take a nuanced stance, acknowledging LLMs’ strengths in PCD tasks while highlighting their limitations in reliably determining the existence of causal relationships.
@@
Full causal graph discovery. Compared with PCD, identifying the full causal graph is a more complicated problem. In a preliminary exploration Long et al. (2023), GPT-3 shows good performance in discovering the causal graph with 3-4 nodes for well-known causal relationships in the medical domain. In more complicated scenarios, the ability of different versions of GPT to discover causal edges Kıcıman et al. (2023) has been validated on the neuropathic pain dataset Tu et al. (2019) with 100 pairs of true/false causal relations. LLM-based discovery (GPT-3.5 and GPT-4) on Arctic sea ice dataset Huang et al. (2021) has comparable or even better performance than representative baselines including NOTEARS Zheng et al. (2018) and DAG-GNN Yu et al. (2019). In Ban et al. (2023), the combination of the causal knowledge generated by LLMs and data-driven methods brings improvement in causal discovery in data from eight different domains with small causal graphs (5$\sim$48 variables and 4$\sim$84 causal relations). However, similar to PCD, LLMs also face many debates about their true ability to discover full graphs Zhou et al. (2024b); Jin et al. (2023b).
@@ 4.2 LLM for Causal Effect Estimation
Although comparatively underexplored, LLMs have also shown impressive performance in causal effect estimation. These works can be mainly categorized into two branches: (1) Causal effect in data: LLMs estimate causal effects within data Lin et al. (2023); Kıcıman et al. (2023) by leveraging their reasoning capabilities and large-scale training data. CLADDER Jin et al. (2023a) benchmarks LLMs for causal effect estimation tasks (e.g., ATE in Rung 2, and ATT, NDE, NIE in Rung 3). Although this task remains challenging, techniques like CoT prompting Jin et al. (2023a) significantly improve performance. (2) Causal effect in models: This branch investigates causal effects involving LLMs themselves, such as the impact of input data, neurons, or learning strategies on predictions Vig et al. (2020); Meng et al. (2022); Stolfo et al. (2022). These studies help understand LLM behavior and support bias elimination Vig et al. (2020), model editing Meng et al. (2022), and robustness analysis Stolfo et al. (2022). For example, Stolfo et al. (2022) explores the causal effect of input (e.g., problem description and math operators) on output solutions in LLM-based mathematical reasoning. In Vig et al. (2020), a causal mediation analysis for gender bias is conducted in language models.
@@ 4.3 LLM for Other Causal Tasks
Experiments Chen et al. (2024a); Jin et al. (2023a); Kıcıman et al. (2023) have shown that there are various other causal tasks that LLMs can bring benefits to. (1) Causal attribution: LLMs show their capability in attribution tasks Kıcıman et al. (2023); Cai et al. (2023) typically in the forms of "why" or "what is the cause" questions. Related tasks also include identifying necessary or sufficient causes Liu et al. (2023); Kıcıman et al. (2023). By embedding human knowledge and cultural common sense, the results show that LLMs have the potential to flexibly address attribution problems in specific domains (such as law and medicine) where conventional methods may fall short Kıcıman et al. (2023). (2) Counterfactual reasoning: Recent studies Kıcıman et al. (2023); Jin et al. (2023a) explore different counterfactual reasoning scenarios, which are often in "what if" questions. While this task is one of the most challenging causal tasks, the improvement in LLMs compared to other methods is noteworthy. (3) Causal explanation: Many recent works explore causal explanations with LLMs Bhattacharjee et al. (2023); Gat et al. (2023); Cai et al. (2023); Gao et al. (2023). Despite ongoing debates regarding LLM’s actual ability for causal reasoning, most empirical studies positively indicate that LLMs serve as effective causal explainers Gao et al. (2023). Such achievement is powered by LLMs’ capability of analyzing language logic and answering questions with natural language.
@@ 4.4 Main Observations and Insights
From the evaluation above and results shown in Table 2 and Table 3, we summarzie the main observations as follows: (1) Model performance: In general, many LLMs exhibit impressive performance in various causal tasks, especially in causal discovery, even with basic prompts. In some cases, their performance can be comparable to or even surpass human-level reasoning Kıcıman et al. (2023). However, as the task difficulty increases from Rung 1$\sim$3, their performance becomes less satisfactory in higher-level complicated causal reasoning tasks Chen et al. (2024a). (2) Enhancement through proper strategies: The performance of LLMs can be significantly enhanced with effective prompting strategies, such as few-shot ICL and CoT. These approaches enable models to leverage contextual information and reasoning processes to achieve better results. Additionally, these models can provide valuable insights through causal explanations. (3) General patterns: While no definitive laws determine model performance universally, certain trends are still observable. For instance, scaling laws suggest that larger models generally perform better, although this is not always that straightforward. These trends provide valuable insights that can guide the future design and development of models. (4) Variability in model effectiveness: There is currently no universally superior LLM or strategy for causal tasks, as their effectiveness can vary significantly depending on the specific scenario. These observations highlight the need for more nuanced and adaptable approaches. (5) Common issues: Current LLMs still struggle with many issues in causal tasks. For example, the answers often lack robustness and are sensitive to changes in prompts Kıcıman et al. (2023); Jin et al. (2023a). Besides, these models frequently default to memorizing and repeating information rather than actual causal reasoning Zečević et al. (2023), which can limit their effectiveness in complex causal scenarios. Besides, LLMs often fail to generate self-consistent causal answers, i.e., the answers from LLMs often conflict with each other. Ongoing debates about whether LLM truly performs causal inference also compel more in-depth analysis.
@@ 5 Discussion and Future Prospects
In general, LLMs offer intriguing perspectives on causal inference, but current research also reveals many limitations, pointing to potential directions for future work that could advance the field Zhang et al. (2023); Kıcıman et al. (2023).
Involving human knowledge: A more comprehensive integration of human knowledge into LLMs can improve causal reasoning, enabling analysis across both general and specialized fields like finance, health, and law Chen et al. (2024b).
Improving data generation: Real-world data often lacks verified causal relations and counterfactuals. LLMs can generate diverse, realistic data with reliable causal relationships, enriching datasets and improving causal reasoning model training.
Addressing hallucinations: In causal reasoning, hallucinations are commonly generated and difficult to detect, leading to misleading causal conclusions. Reducing them is essential to improve the reliability of LLM in causal tasks.
Improving explanation and interactivity: Developing interpretable and instructable LLMs for causal reasoning is crucial. Techniques like fine-tuning, probing, prompt engineering, and optimizing reasoning chains can foster more collaborative and controllable causal inference.
Exploring multimodal causality: Real-world causal scenarios often involve multiple modalities. Recent studies have begun exploring causality across different modalities, such as images Li et al. (2024) and videos Lam et al. (2024). Future research could further investigate these multimodal approaches to enhance causal reasoning.
Developing a unified causal benchmark: There is currently a lack of unified and widely recognized benchmarks for evaluating causal performance for LLMs. Creating a comprehensive benchmark would facilitate LLM assessment.
Advancing causality-specialized models: Most current methods use original LLMs without sufficient focus on causality-centric model designs. There is a significant opportunity for further research and development in specialized causal LLMs to deepen their causal understanding and improve their effectiveness in causal inference.
@@ 6 Limitations
In this survey, it is important to acknowledge certain limitations that shape the scope and focus of our review. First, our analysis is primarily centered on the application of LLMs for causal inference tasks, thereby excluding exploration into how causality is utilized within LLM frameworks themselves. This decision provides a targeted perspective on leveraging LLMs to enhance causal inference methodologies but does not delve into the internal mechanisms or implementations of causal reasoning within these models.
Second, while we comprehensively examine the technical aspects and methodological advancements in using LLMs for causal inference, we do not extensively discuss ethical considerations or potential societal impacts associated with these applications. Ethical dimensions, such as fairness, bias mitigation, and privacy concerns, are critical in the deployment of AI technologies, including LLMs, and warrant dedicated attention and scrutiny in future research and applications. Addressing these limitations ensures a nuanced understanding of the opportunities and challenges in harnessing LLMs for causal inference while also advocating for responsible and ethical AI development and deployment practices.